Circular No. 5658

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA IN M31 James Bryan, Georgetown, TX, reports his discovery of a nova in the

nuclear bulge of M31, found on Technical Pan photographs taken by W. Wren using the McDonald Observatory 0.76-m telescope. The nova is located at α = 0^{h}40^{m}16^{s}.5 ± 0^{s}.5, δ = +40^{o}58′11′′ ± 5′′ (equinox 1950.0). A field star (star 51 of Ciardullo et al. 1987, Ap.J. 318, 520; B = 17.1, V = 16.4) is located 11′′ west-northwest of the nova. The following V magnitudes are available (those by Bryan were from Tech Pan photographs with a 0.40-m telescope): Oct. 26.09 UT, [18.5 (Bryan); Nov. 9.08, 17.2 (Wren); 11.17, 17.2 (Wren); 15.11, 17.2 (Bryan).

A. V. Filippenko, T. Matheson, and L. C. Ho, University of California at Berkeley, report that CCD spectrograms (range 320-1000 nm) obtained on Nov. 18 UT with the Shane 3-m reflector at Lick Observatory reveal that Bryan’s new object is indeed a nova. Intense hydrogen Balmer emission lines are present on a weak continuum, as are fainter lines of Fe II. The spectrum of faint star located nearby (within 1′′-2′′) is superposed on the nova spectrum; thus, part of the weak continuum in the spectrum is probably due to the adjacent star.

GRS 1915+105 AND B1900+14 J. Lochner and L. Whitlock, Goddard Space Flight Center, NASA,

and Universities Space Research Association, communicate: “We have examined the Vela 5B data in the region of the x-ray transient GRS 1915+105 (IAUC 5590, 5619). The Vela 5B data base covers the energy range 3-12 kev from 1969 May to 1979 June. To a sensitivity of 0.1 Crab, we have found no x-ray activity lasting longer than 2.3 days in the region of GRS 1915+105. A recently reported soft γ-ray repeater has been discovered in this region (IAUC 5567, and C. Kouveliotou, Compton Gamma Ray Conf., St. Louis, 1992 Oct.), which may be associated with either the transient or the soft γ-ray repeater B1900+14. To further check the possible association between x-ray transients and soft γ-ray events, we carefully examined the Vela 5B data covering the 1979 March timeframe, in which Mazets et al. (1979, Sov. Astron. Lett. 5, 343) reported events from B1900+14. Good sky coverage allows us to rule out the presence of any x-ray transient event from either GRS 1915+105 or B1900+14 in early 1979.”

1992 November 20 Daniel W. E. Green